OpenStudy (anonymous):

Factor completely 3x3 + 6x2 - 24x. 3x(x - 4)(x + 2) 3x(x + 4)(x - 2) 3x(x - 3)(x + 2) 3x(x + 3)(x + 2)

OpenStudy (anonymous):

@iGreen @_jazmineee @fallenangelorchid @january123 @Lady.Liv1776 @LOUISndGUCCI @sleepyjess @tylermcmullen23

OpenStudy (january123):

I'll pm u

OpenStudy (anonymous):

The answer was not there

OpenStudy (january123):

Classify the expression 7x3 + 2x2 – 9x as a monomial, binomial, trinomial or polynomial. Answer monomial binomial polynomial trinomial <==ANSWER Identify the degree of the polynomial 8x3y2 – 10xy + 4x2y2 + 3. Answer 2 4 5 <==ANSWER, first term is 3 + 2 3 Identify the polynomial written in ascending order. Answer 3x3 – 5x2 + 4x – 7 –5x2 + 4x + 3x3 – 7 –7 + 4x – 5x2 + 3x3 <==ANSWER 4x + 3x3 – 7 – 5x2 Simplify (4x2 + 12x – 9) + (7x2 – x + 3) Answer 11x2 + 12x – 6 11x2 + 13x + 6 11x2 + 11x – 6 <==ANSWER 11x2 + 11x + 6 Simplify (5x3 – x + 14) – (3x2 – 9x + 4) Answer 5x3 – 3x2 + 8x + 10 <==ANSWER 2x2 – 10x + 18 5x3 – 3x2 – 10x + 18 2x2 + 8x + 10 Simplify 4x2(2x2 – 8x + 6) Answer 8x4 – 32x3 + 24x2 <==ANSWER 6 – 32x + 24x2 8x2 – 32x + 24 6x2 – 32x + 24 Which of the following represents the area of a rectangle whose length is x – 7 and whose width is x + 12? Answer x2 – 5x – 84 <==ANSWER x2 + 5x – 84 x2 – 84 x2 – 84x + 5s . Simplify (x + 4)(x2 – 6x + 3) Answer x3 – 6x2 – 17x + 12 x3 – 10x2 – 27x + 12 x3 – 14x2 + 3x + 12 x3 – 2x2 – 21x + 12 <==ANSWER . Factor completely 4x4 – 20x3 + 12x2 Answer 4x2(x2 – 5x + 3) <==ANSWER 4x2(x – 3)(x – 1) 4(x4 – 5x3 + 3x2) 4(x – 3)(x – 1) Factor completely 81x2 – 100 Answer (9x – 10)(9x + 10) <==ANSWER (10x – 9)(10x + 9) (9x – 10)(9x – 10) (10x – 9)(10x – 9) Factor completely 100x2 + 60x + 9 Answer None of the above <==ANSWER, should be (10x + 3)² (10x – 3)(10x + 3) (3x – 10)2 (10x – 3)2 Which of the following is a factor of x2 – 21x + 110? Answer x + 11 x – 5 None of the above x – 10 <==ANSWER Which is a factor of 2x2+ 8x – 90 ? Answer x + 9 <==ANSWER x + 2 x – 4 x – 15 Factor completely x2 – 4x – 12 Answer (x – 2)(x – 6) prime (x + 2)(x – 6) <==ANSWER (x – 2)(x + 6) Which is a factor of 3x2y2 + 6x2+ 12y2 + 24 ? 3(x2 + 4)(y2 + 2) Answer y2 + 6 6 x2 + 4 <==ANSWER x2 + 2 Which of the following is a factor of 15x2 – 14x – 8? (15x - 20)(15x + 6) ------------- ------------ = (3x - 4)(5x + 2) . . . 5 . . . . . . 3 Answer 5x – 2 None of the above 5x + 16 3x – 4 <==ANSWER Which is a factor of 24x2– 44x + 12 ? (24x - 36)(24x - 8) ------------- ------------ = 4(2x - 3)(3x - 1) . . . 12 . . . . . 8 Answer 3x + 3 3x – 2 3 2x – 3 <==ANSWER Which of the following is a factor of 512x3 – 125? Answer 8x + 5 64x2 + 80x + 25 64x2 + 40x + 25 <==ANSWER None of the above Which of the following is a factor of 32x3 + 4? Answer 2x – 1 None of the above 4x2 – 4x + 1 <==ANSWER 4x2 – 2x + 1 Which is not a factor of 18x3 – 123x2 – 21x? 3x(6x2 - 41x - 7) = 3x(x - 7)(6x + 1) Answer (6x + 1) (x – 7) 3x (3x – 2) <==ANSWER I hope that helps!! :-)

OpenStudy (anonymous):

Thanks but not really